Circular No. 6788

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971214 N. Itoh, Y. Nakada, and T. Soyano, Kiso Observatory, University of

Tokyo, report: “The GRB error circle was surveyed done to I = 20 and R = 21 with the Kiso 1.05-m Schmidt telescope on Dec. 15.8 UT. The R-band image shows no new objects brighter than the limiting magnitude of the second Palomar Sky Survey. The I-band image from three 10-min exposures revealed one faint object 54′′ northeast of the BeppoSAX position, at α = 11^{h}56^{m}25^{s}.2, δ = +65^{o}11′57′′.9 (equinox 2000.0; ± 6′′), I = 19.7 ± 0.5.”

J. Halpern, J. Thorstensen, D. Helfand, E. Costa, and the BeppoSAX team have detected a fading optical transient within the error circle of GRB 971214 (IAUC 6787) in a series of I-band images taken on the MDM Observatory 2.4-m telescope on Kitt Peak. The position is α = 11^{h}56^{m}26^{s}.40, δ = +65^{o}12′00′′.5 (equinox 2000.0). The I magnitude was 21.2 ± 0.3 on Dec. 15.47 UT, and near the detection limit of 22.6 on Dec. 16.47. Images

will be made available at http://www.astro.columbia.edu/~cba/mdm/.

XTE J0053–724 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association (USRA); F. E. Marshall, GSFC; J. C. Lochner, GSFC and USRA; and M. Ozaki and Y. Ueda, Institute of Space and Astronautical Science, report: “After the apparent detection of SMC X-3 with RXTE (IAUC 6777), we performed ASCA observations on Dec. 12. While no source was found at the location of SMC X-3 itself, a source was detected ∼ 8′ away, at α = 0^{h}53^{m}53^{s}, δ = −72^{o}26′42′′ (equinox 2000.0; ± 1′.2, 90percent confidence) that exhibited modulation at the 92-s period seen with RXTE, confirming that this is the same source. The mean countrate (per GIS detector) of 0.08 counts/s corresponds to a 2–10-keV flux of 1.45 × 10^{−11} erg s^{−1} cm^{−2}. A search of archival ROSAT images reveals a variable source in this error box located at position end figures 55^{s}, 48′′ (± 5′′).”

Lochner and L. A. Whitlock, GSFC and USRA; and N. Brandt, Pennsylvania State Universtity, report: “In a preliminary analysis of the RXTE PCA observations of XTE J0053–724 from Nov. 25–Dec. 11, the spectra have been fit with a cutoff power law (with photon indices typically 0.5–0.7 and energy cutoffs in the range 11–15 kev) show no evidence for intrinsic x-ray absorption. Flux measurements in the range 2–10 keV are as follows (units of 10^{−11} erg s^{−1} cm^{−2}): Nov. 25.46 UT, 6.72; 27.45, 6.74; 29.25, 6.51; Dec. 2.18, 6.06; 2.35, 5.97; 5.18, 5.30; 8.06, 4.82; 11.85, 4.12. The average pulsation period measured from these observations is 91.13 s.”

1997 December 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green